# Neural Combinatorial Optimization Algorithms for Solving Vehicle Routing Problems: A Comprehensive Survey with Perspectives

### v2

@@ Abstract
Although several surveys on Neural Combinatorial Optimization (NCO) solvers specifically designed to solve Vehicle Routing Problems (VRPs) have been conducted. These existing surveys did not cover the state-of-the-art (SOTA) NCO solvers emerged recently. More importantly, to provide a comprehensive taxonomy of NCO solvers with up-to-date coverage, based on our thorough review of relevant publications and preprints, we divide all NCO solvers into four distinct categories, namely Learning to Construct, Learning to Improve, Learning to Predict-Once, and Learning to Predict-Multiplicity solvers. Subsequently, we present the inadequacies of the SOTA solvers, including poor generalization, incapability to solve large-scale VRPs, inability to address most types of VRP variants simultaneously, and difficulty in comparing these NCO solvers with the conventional Operations Research algorithms. Simultaneously, we propose promising and viable directions to overcome these inadequacies. In addition, we compare the performance of representative NCO solvers from the Reinforcement, Supervised, and Unsupervised Learning paradigms across both small- and large-scale VRPs. Finally, following the proposed taxonomy, we provide an accompanying web page as a live repository for NCO solvers. Through this survey and the live repository, we hope to make the research community of NCO solvers for VRPs more thriving.
@@ I Introduction
Combinatorial Optimization Problem (COP) is a vital branch of mathematical optimization, dedicated to seeking optimal solutions for problems within the discrete space [1, 2, 3]. Among various COPs, Vehicle Routing Problems (VRPs) have extensive practical applications across diverse domains, e.g., communication and transportation [4, 5, 6]. Generally speaking, Operations Research (OR) algorithms for solving VRPs can be broadly divided into three categories, namely exact, approximation, and heuristic algorithms [7]. Specifically, exact algorithms typically employ the divide-and-conquer (D&C) manner to solve VRPs and obtain optimal solutions. These algorithms mainly encompass Branch and Bound [8] and Dynamic Programming (DP) [9]. Approximation algorithms, aimed at securing solutions of assured quality, include Relaxation Algorithm [10] and Linear Programming [11]. Heuristic algorithms represent a class of methods that utilize the predefined heuristic rules to efficiently explore solutions in a given timeframe, developed through continuous trial-and-error. The most prominent heuristic algorithm for solving VRPs is Lin-Kernighan-Helsgaun 3 (LKH3) [12]. Additionally, it is noteworthy that approximation and heuristic algorithms may have the potential to converge to the optimal solutions given sufficient time, yet such convergence is not guaranteed [13, 7].
@@
However, these three types of conventional OR algorithms lack the ability to derive insights from historical VRP instances, leading to significant computing overhead, especially for medium- or large-scale VRPs [14]. This is demonstrated by the performance of LKH3 algorithm in solving a 100-node Capacity Vehicle Routing Problem (CVRP) instance, requiring a noteworthy 12 hours to attain the optimal solution. This shortcoming constrains the applications of OR algorithms in time-sensitive scenarios, such as on-call routing [15] and ride-hailing service [16]. To alleviate the expensive computing overhead, many Neural Combinatorial Optimization (NCO) solvers utilizing Deep Learning (DL) have recently emerged [17, 18, 19, 20, 21, 22], as shown in Fig. 1. These NCO solvers possess the ability to learn from extensive historical instances and efficiently search for (sub-)optimal solutions, when dealing with instances exhibiting the same distribution characteristics. Moreover, by leveraging Graphics Processing Units (GPUs), these NCO solvers demonstrate efficiency in concurrently handling batches of instances, thereby further enhancing the speed of problem-solving [23, 24, 25, 26]. For example, for a 100-node CVRP instance, an NCO solver named POMO [27] requires only one minute to find the sub-optimal solution with a gap value of less than 1%. Here, the gap denotes the relative difference in the “average length” between POMO and LKH3.
In recent three years, a number of surveys [23, 28, 7, 17, 29, 30, 31, 32, 33, 18, 34] have been conducted to introduce and summarize NCO solvers specifically designed for VRPs. Nonetheless, these existing surveys have various limitations. First, the taxonomies proposed in these surveys [23, 28, 7, 17, 29] do not facilitate an appropriate category for all NCO solvers. Secondly, a number of surveys [7, 17, 29, 30, 31, 32, 18, 34, 33] do not adequately cover current NCO solvers’ inadequacies and associated efforts. Thirdly, a few surveys [30, 31, 32] exclusively focus on NCO solvers utilizing Reinforcement Learning (RL) paradigm, overlooking those based on Supervised Learning (SL) and Unsupervised Learning (UL). Finally, many surveys [23, 7, 17, 29, 30, 31] focus on multiple COPs, without providing a detailed introduction to VRPs. In the following paragraphs, we elaborate on these four limitations of existing surveys and introduce the advantages of this survey.
@@
Category Representatives Description
L2C Ptr-Net [25]; AM [26]; POMO [27]; NAR4TSP[38] Using NNs to construct solutions from scratch
L2I NeuRewriter [39]; MT [40]; N2OPT [41]; NeuOPT [42] Using NNs to iteratively improve complete solutions
L2P-O GCN [37]; NeuroLKH [43]; DPDP [44]; DIFUSCO [45] Using NNs to assist OR algorithms once
L2P-M GAVE [46]; VSR-LKH [47]; MOCO [48]; EOH [49] Using NNs to assist OR algorithms Multiplicity
@@
In this survey, we propose a novel taxonomy aimed at appropriately and exclusively classifying all NCO solvers. As shown in Table I, we categorize these solvers into four categories, namely Learning to Construct (L2C), Learning to Improve (L2I), Learning to Predict-Once (L2P-O), and Learning to Predict-Multiplicity (L2P-M) solvers. L2C solvers represent the pioneering category of NCO solvers, drawing inspiration from the analogy between VRPs and the machine translation task. Specifically, L2C solvers use NNs to sequentially select nodes that have not been visited yet (resembling the process of translating words in the machine translation task) and add them to partial solutions [26, 50, 51]. For example, Vinyals et al. [25] were the pioneers who employed SL to train a Recurrent Neural Network (RNN) for sequentially constructing solutions for Traveling Salesman Problem (TSP). Subsequently, inspired by the notion of iteratively ruining and repairing in heuristic algorithms [52], certain L2I solvers [6, 39, 53, 54] introduce NNs to emulate this process, aiming to derive the (sub-)optimal solutions. For example, Chen and Tian [39] selectively deconstruct the local components of complete solutions and repair them through a learned policy, enhancing the quality of current complete solutions. Meanwhile, certain L2P-O solvers [37, 44] and L2P-M solvers [47, 46] diverge from emulating the iteratively improving process. Instead, these solvers focus on integrating the predicted key information with OR algorithms to enhance solution quality and/or reduce search overhead. Furthermore, L2P-O solvers predict information only once throughout the entire search process of OR algorithms, whereas L2P-M ones require the facilitation of OR algorithms during the whole search process, making predictions and decisions multiple times. For example, to expedite the search process of DP, Kool et al. [44] proposed an L2P-O solver named Deep Policy Dynamic Programming (DPDP). This solver predicts promising edges for each TSP instance before the initiation of the DP search. The solution is then exclusively constructed using these identified promising edges during the subsequent DP search. In line with the objective of [44] to expedite the search process of OR algorithms, an L2P-M solver named VSR-LKH [47] exploits RL to inform decisions at each search step of the LKH algorithm [55]. In conclusion, the taxonomy proposed in this paper provides a structured framework to comprehensively present the diversity of NCO solvers, ensuring that all solvers can be exclusively categorized into one of the four categories defined in our taxonomy. In Sections III to V, we discuss the reasonableness and exclusivity of the proposed taxonomy, providing comprehensive overviews of the relevant studies corresponding to each solver category.
@@
The second limitation arises from the lack of comprehensive reporting on the inadequacies of state-of-the-art (SOTA) NCO solvers in prior surveys [7, 17, 29, 30, 31, 32, 18]. Instead, they only introduce the relatively outdated inadequacies of NCO solvers, failing to provide researchers and practitioners with insights into the latest developments in NCO solvers. For example, certain surveys identify the performance of NCO solvers on small-scale VRPs as a potential inadequacy. However, recent NCO solvers (e.g., SGBS [56]) have effectively addressed this inadequacy. To identify the open inadequacies of SOTA NCO solvers, we conduct a systematic survey and simultaneously introduce on-going efforts to address four open inadequacies. Specifically, the first inadequacy is that existing NCO solvers are only capable of solving VRPs that have the same instance distribution [57, 58, 59, 60]. For example, NCO solvers trained on a type of specific instance distribution (e.g., the uniform distribution) might not perform well when being applied onto instances from other distribution patterns (e.g., the cluster distribution) [58]. In addition, existing NCO solvers face inadequacies in addressing large-scale VRPs (e.g., the 10,000-node TSP) in real-time [61, 62, 63]. Moreover, existing NCO solvers exhibit limited performance on various VRP variants (e.g., Asymmetric Traveling Salesman Problem [64]), necessitating the development of specific NNs for each VRP variant. Finally, the issue of unfair comparisons emerges as a substantial inadequacy in evaluating NCO solvers against the conventional OR algorithms, resulting in a constrained acceptance of these NCO solvers within the OR community [65, 66]. More importantly, apart from summarizing these inadequacies, we also suggest promising research directions to address them (see Section VI for more details).
@@
The third limitation stems from the fact that certain surveys [30, 31, 32, 67] exclusively focus on NCO solvers utilizing RL, overlooking those based on SL and UL. Consequently, these surveys do not provide a comprehensive performance comparison across all solvers, which is a non-negligible issue due to the fact that certain SOTA SL and UL-based solvers [68, 45] have demonstrated high-level performance on large-scale VRPs. In contrast, our survey stands out by introducing NCO solvers based on all learning paradigms, namely SL, UL, and RL. For example, we introduce the SL-based solver LEHD [36] and the UL-based solver UTSP [68], both of which are specifically designed to solve large-scale VRPs by reducing the number of parameters (see Section VI-B for more details). By employing this comprehensive survey methodology, this survey paper provides a thorough performance comparison among solvers across varying scales of VRPs.
@@
Finally, a number of existing surveys [23, 7, 17, 29, 30, 31] have delved into a spectrum of COPs, such as Job Shop Scheduling Problem (JSP) [69, 70, 71] and Knapsack Problem (KP) [72, 73, 74]. Consequently, these surveys are unable to provide a comprehensive and detailed introduction to VRPs. In contrast, this survey paper takes a focused approach to comprehensively introduce the details of NCO solvers tailored for VRPs, the most researched COPs in the NCO domain [75]. For example, in Section III-D, we elaborate on data augmentation methods specially designed for VRPs, which are often overlooked in other surveys. Moreover, another reason for the exclusive focus of this survey on VRPs is that most NCO studies initially focus on VRPs before extending their applications to other COPs. For example, Kool et al. [26] pioneered the integration of Attention Model (AM) into NCO to solve VRPs, which is expanded upon by subsequent studies [71] and [73] to tackle JSPs and KPs, respectively. Hence, it is justified for us to purposely focus on VRPs in this survey.
@@
  • This paper outlines the inadequacies of existing NCO solvers, including generalization, large-scale and multi-constrained VRPs, and fair comparisons with OR algorithms, while suggesting promising and viable solutions.
@@ II Preliminary
Section II-A first introduces the definitions of TSP and CVRP. Subsequently, Section II-B presents three learning paradigms of DL, namely SL, UL and RL.
@@ II-A Definitions of TSP and CVRP
VRPTW Node visits must adhere to specified time windows. Vehicles can wait in the event of early arrival, but late arrival is not permissible. [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104]
MOVRP In addition to minimizing tour length, additional objectives include minimizing traveling time, etc. [105, 106, 107, 108, 109, 110, 111, 112, 113, 114]
DVRP Dynamically scheduling vehicles to fulfill online-generated orders across multiple locations. [115, 116, 117, 118, 119, 120, 121, 122]
@@ II-B1 Supervised Learning (SL)
SL is a paradigm in DL where the model undergoes training on a labeled dataset, containing input-output pairs. The objective of the model is to learn a function that produces predicted outputs closely matching the corresponding output labels, when given a set of input data. In the field of NCO, Vinyals et al. [25] were the first to use a supervised loss function for training an RNN. The loss of a given solution is computed as follows:

### v3

@@ Abstract
Although several surveys on Neural Combinatorial Optimization (NCO) solvers specifically designed to solve Vehicle Routing Problems (VRPs) have been conducted, they did not cover the state-of-the-art (SOTA) NCO solvers emerged recently. More importantly, to establish a comprehensive and up-to-date taxonomy of NCO solvers, we systematically review relevant publications and preprints, categorizing them into four distinct types, namely Learning to Construct, Learning to Improve, Learning to Predict-Once, and Learning to Predict-Multiplicity solvers. Subsequently, we present the inadequacies of the SOTA solvers, including poor generalization, incapability to solve large-scale VRPs, inability to address most types of VRP variants simultaneously, and difficulty in comparing these NCO solvers with the conventional Operations Research algorithms. Simultaneously, we discuss on-going efforts, identify open inadequacies, as well as propose promising and viable directions to overcome these inadequacies. Notably, existing efforts focus on only one or two of these inadequacies, with none attempting to address all of them concurrently. In addition, we compare the performance of representative NCO solvers from the Reinforcement, Supervised, and Unsupervised Learning paradigms across VRPs of varying scales. Finally, following the proposed taxonomy, we provide an accompanying web page as a live repository for NCO solvers. Through this survey and the live repository, we aim to foster further advancements in the NCO community.
@@ I Introduction
Combinatorial Optimization Problem (COP) is a vital branch of mathematical optimization, dedicated to seeking optimal solutions for problems within the discrete space [1, 2, 3]. Among various COPs, Vehicle Routing Problems (VRPs) have extensive practical applications across diverse domains, e.g., communication and transportation [4, 5, 6, 7]. Generally speaking, Operations Research (OR) algorithms for solving VRPs can be broadly divided into three categories, namely exact, approximation, and heuristic algorithms [8]. Specifically, exact algorithms typically employ the divide-and-conquer (D&C) manner to solve VRPs and obtain optimal solutions. These algorithms mainly encompass Branch and Bound [9] and Dynamic Programming (DP) [10]. Approximation algorithms, aimed at securing solutions of assured quality, include Relaxation Algorithm [11] and Linear Programming [12]. Heuristic algorithms represent a class of methods that utilize the predefined heuristic rules to efficiently explore solutions in a given timeframe, developed through continuous trial-and-error. The most prominent heuristic algorithm for solving VRPs is Lin-Kernighan-Helsgaun 3 (LKH3) [13]. Additionally, it is noteworthy that approximation and heuristic algorithms may have the potential to converge to the optimal solutions given sufficient time, yet such convergence is not guaranteed [14, 8].
@@
However, these three types of conventional OR algorithms lack the ability to derive insights from historical VRP instances, leading to significant computing overhead, especially for medium- or large-scale VRPs [15, 16]. This is demonstrated by the performance of LKH3 algorithm in solving a 100-node Capacitied Vehicle Routing Problem (CVRP) instance, requiring a noteworthy 12 hours to attain the optimal solution [17]. This shortcoming constrains the applications of OR algorithms in time-sensitive scenarios, such as on-call routing [18] and ride-hailing service [19]. To alleviate the expensive computing overhead, many Neural Combinatorial Optimization (NCO) solvers utilizing Deep Learning (DL) have recently emerged [20, 21, 22, 23, 24, 25, 26], as shown in Fig. 1. These NCO solvers possess the ability to learn from extensive historical instances and efficiently search for (sub-)optimal solutions, when dealing with instances exhibiting the same distribution characteristics. Moreover, by leveraging GPUs, these NCO solvers demonstrate efficiency in concurrently handling batches of instances, thereby further enhancing the speed of problem-solving [27, 28, 29, 30]. For example, for a 100-node CVRP instance, an NCO solver named POMO [17] requires only one minute to find the sub-optimal solution with a gap value of less than 1%. Here, the gap denotes the relative difference in the “average length” between POMO and LKH3.
In the past three years, a number of surveys [27, 31, 8, 20, 32, 33, 34, 35, 36, 21, 37] have been conducted to introduce and summarize NCO solvers specifically designed for VRPs. Nonetheless, these existing surveys have various limitations. First, the taxonomies proposed in these surveys [27, 31, 8, 20, 32] do not facilitate an appropriate category for all NCO solvers. Secondly, a number of surveys [8, 20, 32, 33, 34, 35, 21, 37, 36] do not adequately cover current NCO solvers’ inadequacies and associated efforts. Thirdly, a few surveys [33, 34, 35] exclusively focus on NCO solvers utilizing the Reinforcement Learning (RL) paradigm, overlooking those based on Supervised Learning (SL) and Unsupervised Learning (UL). Consequently, these surveys do not provide a comprehensive overview of state-of-the-art (SOTA) NCO solvers. Finally, many surveys [27, 8, 20, 32, 33, 34] focus on multiple COPs, without providing a detailed introduction to VRPs. All these necessitate a more up-to-date survey to include the SOTA advancements in this important research field. In the following paragraphs, we elaborate on these four limitations of existing surveys and introduce the advantages of this survey.
@@
Category Description Representatives
L2C Using NNs to construct solutions from scratch Ptr-Net [29]; AM [30]; POMO [17]; NAR4TSP[41]
L2I Using NNs to iteratively improve complete solutions NeuRewriter [42]; MT [43]; N2OPT [44]; NeuOPT [45]
L2P-O Using NNs to assist OR algorithms once GCN [40]; NeuroLKH [46]; DPDP [47]; DIFUSCO [48]
L2P-M Using NNs to assist OR algorithms multiplicity GAVE [49]; VSR-LKH [50]; MOCO [51]; EOH [52]
@@
In this survey, we propose a novel taxonomy aimed at appropriately and exclusively classifying all NCO solvers. As shown in Table I, we categorize these solvers into four categories, namely Learning to Construct (L2C), Learning to Improve (L2I), Learning to Predict-Once (L2P-O), and Learning to Predict-Multiplicity (L2P-M) solvers. L2C solvers represent the pioneering category of NCO solvers, drawing inspiration from the analogy between VRPs and the machine translation task. Specifically, L2C solvers use NNs to sequentially select nodes that yet to be visited (resembling the process of translating words in the machine translation task) and add them to partial solutions [30, 53, 54]. For example, Vinyals et al. [29] were the pioneers who employed SL to train a Recurrent Neural Network (RNN) for sequentially constructing solutions for Traveling Salesman Problem (TSP). Subsequently, inspired by the notion of iteratively ruining and repairing in heuristic algorithms [55], certain L2I solvers [6, 42, 56, 57] introduce NNs to emulate this process, aiming to derive the (sub-)optimal solutions. For example, Chen and Tian [42] selectively deconstructed the local components of complete solutions and repaired them through a learned policy, enhancing the quality of current complete solutions. Meanwhile, certain L2P-O solvers [40, 47] and L2P-M solvers [50, 49] diverge from emulating the iteratively improving process. Instead, these solvers focus on integrating the predicted key information with OR algorithms to enhance solution quality and/or reduce search overhead. Furthermore, L2P-O solvers predict information only once throughout the entire search process of OR algorithms, whereas L2P-M ones require the facilitation of OR algorithms during the whole search process, making predictions and decisions multiple times. For example, to expedite the search process of DP, Kool et al. [47] proposed an L2P-O solver named Deep Policy Dynamic Programming (DPDP). This solver predicts promising edges for each TSP instance before the initiation of the DP search. The solution is then exclusively constructed using these identified promising edges during the subsequent DP search. In line with the objective of [47] to expedite the search process of OR algorithms, an L2P-M solver named VSR-LKH [50] exploits RL to inform decisions at each search step of the LKH algorithm [58]. In conclusion, the taxonomy proposed in this paper provides a structured framework to comprehensively present the diversity of NCO solvers, ensuring that all solvers can be exclusively categorized into one of the four categories defined in our taxonomy. In Sections III to V, we discuss the reasonableness and exclusivity of the proposed taxonomy, providing comprehensive overviews of the relevant studies corresponding to each solver category.
@@
The second limitation arises from the lack of comprehensive reporting on the inadequacies of SOTA NCO solvers in prior surveys [8, 20, 32, 33, 34, 35, 21]. Unfortunately, they only introduce the relatively outdated inadequacies of NCO solvers, failing to provide researchers and practitioners with insights into the latest developments in NCO solvers. For example, certain surveys identify the performance of NCO solvers on small- or medium-scale VRPs as a potential inadequacy. However, recent NCO solvers (e.g., GELD [59]) have effectively addressed this inadequacy, achieving performance on par with exact algorithms. Again, a more up-to-date survey is required to reflect these recent advancements. To identify the open inadequacies of SOTA NCO solvers, we conduct a systematic survey and simultaneously introduce on-going efforts to address four open inadequacies. Specifically, the first inadequacy is that existing NCO solvers are only capable of solving VRPs that have the same instance distribution [60, 61, 62, 63]. For example, NCO solvers trained on a type of specific instance distribution (e.g., the uniform distribution) might not perform well when being applied onto instances from other distribution patterns (e.g., the cluster distribution) [61]. In addition, existing NCO solvers face inadequacies in addressing large-scale VRPs (e.g., the 10,000-node TSP) in real-time [64, 65, 66]. Moreover, existing NCO solvers exhibit limited performance on various VRP variants (e.g., Asymmetric Traveling Salesman Problem [67]), necessitating the development of specific NNs for each VRP variant. Finally, the issue of unfair comparisons emerges as a substantial inadequacy in evaluating NCO solvers against the conventional OR algorithms, resulting in a constrained acceptance of these NCO solvers within the OR community [68, 69]. To better solve these inadequacies, recent studies [59, 70, 71] have explored various approaches and proposed corresponding solvers, which have been reviewed in this survey. More importantly, apart from introducing on-going efforts to address these inadequacies, we also identify open inadequacies and suggest promising research directions to address them (see Section VI for more details). It is worth highlighting that while certain prior studies (e.g., GELD [59]) have attempted to tackle one or two of these inadequacies, none simultaneously addresses three or even all four inadequacies, and the overall performance of many NCO solvers still leaves room for improvement (see Section VI-C for more details).
@@
The third limitation stems from the fact that certain surveys [33, 34, 35, 72] exclusively focus on NCO solvers utilizing RL, overlooking those based on SL and UL. Consequently, these surveys do not provide a comprehensive performance comparison across all solvers, which is a non-negligible issue due to the fact that certain SOTA SL- and UL-based solvers [73, 48] have demonstrated high-level performance on large-scale VRPs. In contrast, our survey stands out by introducing NCO solvers based on all learning paradigms, namely SL, UL, and RL. For example, we introduce the SL-based solver LEHD [39] and the UL-based solver UTSP [73], both of which were specifically designed to solve large-scale VRPs by reducing the number of parameters (see Section VI-B for more details). By employing this comprehensive survey methodology, this survey paper provides a thorough performance comparison among solvers across varying scales of VRPs.
@@
Finally, a number of existing surveys [27, 8, 20, 32, 33, 34] have delved into a spectrum of COPs, such as Job Shop Scheduling Problem (JSP) [74, 75, 76, 77] and Knapsack Problem (KP) [78, 79, 80]. Consequently, these surveys are unable to provide a comprehensive and detailed introduction to VRPs. In contrast, this survey paper takes a focused approach to comprehensively introduce the details of NCO solvers tailored for VRPs, the most researched COPs in the NCO domain [81]. For example, in Section III-D, we elaborate on data augmentation methods specially designed for VRPs, which are often overlooked in other surveys. Moreover, another reason for the exclusive focus of this survey on VRPs is that most NCO studies initially focused on VRPs before extending their applications to other COPs. For example, Kool et al. [30] pioneered the integration of Attention Model (AM) into NCO to solve VRPs, which is expanded upon by subsequent studies [76] and [79] to tackle JSPs and KPs, respectively. Hence, it is justified for us to purposely focus on VRPs in this survey.
@@
  • This paper outlines the inadequacies of existing NCO solvers, including generalization, large-scale and multi-constrained VRPs, and fair comparisons with OR algorithms, while introducing on-going efforts and suggesting promising directions to better address these inadequacies.
@@ II Preliminary
Section II-A first introduces the definitions of TSP and CVRP. Subsequently, Section II-B presents three learning paradigms of DL, namely SL, UL, and RL.
@@ II-A Definitions of TSP and CVRP
VRPTW Node visits must adhere to specified time windows. Vehicles can wait in the event of early arrival, but late arrival is not permissible. [99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110]
MOVRP In addition to minimizing tour length, additional objectives include minimizing traveling time, etc. [111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122]
DVRP Dynamically scheduling vehicles to achieve online-generated orders across multiple locations. [123, 124, 125, 126, 127, 128, 129, 130]
@@ II-B1 Supervised Learning (SL)
SL is a paradigm in DL where the model undergoes training on a labeled dataset, containing input-output pairs. The objective of the model is to learn a function that produces predicted outputs closely matching the corresponding output labels, when given a set of input data. In the field of NCO, Vinyals et al. [29] were the first to use a supervised loss function for training an RNN. The loss is computed as follows: